OpenStudy (anonymous):

a person sitting on the top of a building is dropping balls at regular interval of 1 sec .find position of the 3rd ,4th, and 5th ball when 6th ball is being drop.

OpenStudy (anonymous):

use h=1/2 *g *t2 and then successively put for t as 1 , 2,3,4,5 ...

OpenStudy (anonymous):

matricked gives the correct formula to use; the height will be determined by putting in the correct times ie 3 seconds, 2 seconds and 1 second

OpenStudy (anonymous):

Dn = U +a(n-1/2) n= time substitude and get the answer

OpenStudy (anonymous):

hlp me out/